# Domain-Specific Translation with Open-Source Large Language Models: Resource-Oriented Analysis

### v2

@@ 5 Conclusions
After a few years of research on leveraging open-source LLMs for direct translation among other tasks, several businesses have been investigating the feasibility of integrating such general-purpose models into their specialized translation workflows. Among the main advantages of employing the in-context learning feature of LLMs is maintaining consistency, domain terminology and style throughout the entire document Dettmers et al., (2023); Wang et al., (2023). This can be achieved by feeding an LLM with similar translation pairs, preapproved terms, or other useful data at inference time, hoping that the LLM’s output will be influenced by the provided linguistic data. Moreover, several businesses and organizations have resorted to open-source LLMs for data privacy and security reasons. Nevertheless, deploying open-source LLMs into production poses efficiency challenges, including infrastructure requirements and latency at translation time. This raises the question of whether conventional encoder-decoder MT models remain relevant, particularly for domain-specific translation.
As elaborated in Section 3, we evaluated a range of LLMs in the medical domain. Moreover, we fine-tuned selected models for zero-shot and one-shot translation. Our experiments covered both the multilingual encoder-decoder MT model, NLLB-200 3.3, and several autoregressive decoder-only LLMs, such as Gemma, Llama, and Mistral. Our retrieval-augmented translation used one-shot in-context examples based on semantic similarly with the new source (fuzzy match), which achieves better results than random selection (Agrawal et al.,, 2023; Moslem et al., 2023a, ). In the future, we would like to compare retrieval-augmented context (as in this work) with direct context from previous and next sentences in the document.
@@ 6 Limitations
Since all the data used in this work has been publicly available for years, there is a potential for data contamination. Many open-source LLMs do not disclose the specific datasets used for training, so we cannot be certain whether our test datasets were part of the training data for any of these models. However, we conducted multiple experiments for each model, including sentence-level translation (zero-shot), retrieval-augmented translation with fuzzy matches (one-shot), and fine-tuning of some models. As such, we consider our results to provide a reliable indication of the translation quality that can be expected from these LLMs in similar setups. It is important to note that, as this paper demonstrates, results may vary across different languages or domains.
@@ A.1 Filtering
Filtering
Dataset Raw Rule-Based Semantic
ELRC_2922 4,365 4,241 4,241
ELRC_2923 502 486 486
ELRC_3382 6,593 6,547 6,547
ELRC-antibiotic 1,072 1,051 1,051
ELRC-EUROPARL_covid 789 773 773
ELRC-presscorner_covid 11,361 11,296 11,296
ELRC-wikipedia_health 4,366 4,269 4,269
EMEA 1,092,568 313,240 313,240
Tico-19 3,071 2,971 2,971
Total 1,156,317 376,216 376,216
@@
Filtering
Dataset Raw Rule-Based Semantic
ELRC_2922 7,292 7,197 7,197
ELRC_2923 560 545 545
ELRC_3382 3,513 3,479 3,479
ELRC-antibiotic 839 821 821
ELRC-EUROPARL_covid 864 843 843
ELRC-presscorner_covid 6,319 6,263 6,263
ELRC-wikipedia_health 7,293 7,223 7,223
EMEA 1,082,144 310,619 310,619
SciELO 3,084,830 2,732,045 2,732,045
Tico-19 3,071 3,069 3,069
Total 4,196,725 3,072,104 3,072,104
@@
Filtering
Domain Dataset Raw Rule-Based Semantic
Medical ELRC-wikipedia_health 608 594 593
Tico-19 3,071 3,058 3,056
Total 3,679 3,652 3,649
@@
Generic GlobalVoices 32,307 29,633 29,623
Wikimedia 16,289 13,834 13,829
Total 48,596 43,467 43,452
@@ B.2.2 NLLB
B.3 Dataset Packing
This section examines the performances of the models with and without packing. Packing dataset, initially introduced by (Raffel et al.,, 2020), helps increase the training efficiency by enabling fine-tuning on more sequences. In this method, multiple short examples are packed in the same input sequence. Packing tokens into the maximum sequence length helps improve training efficiency. While packing is primarily intended to improve the training throughput, our experiments demonstrate it also improves model performance. In our experiments, we use the HuggingFace’s SFTTrainer, which supports example packing.
@@ C.1 CTranslate2 vs vLLM
C.2 Quantization
Moreover, we compared generation with Llama-3 70B with and without Activation-aware Weight Quantization (AWQ) using vLLM. As Table 14 shows, results without quantization are slightly better. All the results in the main paper for LLMs are for non-quantized models, except Llama-3.1 405B and DeepSeek V3 685B, due to commutating constrains. Hence, we tested Llama-3 70B with and without quantization and reported the results in this section to understand the effect of AWQ on the translation quality.
Lang Context Quant. BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR zero-shot None 39.74 62.51 53.36 66.75
AWQ 39.02 61.97 54.69 64.97
one-shot None 51.93 70.57 43.03 76.76
AWQ 51.83 70.43 43.00 76.19
EN-PT zero-shot None 43.86 67.35 43.95 84.05
AWQ 43.12 66.59 45.58 82.12
one-shot None 45.95 68.82 41.68 86.58
AWQ 45.59 68.57 42.12 86.10
@@ C.2 Quantization
Table 15: Evaluation of the translation performance of NLLB-200 3.3B and Mistral 7B, both without quantization and with lower-precision quantization formats (float16 and int8) using CTranslate2. The results indicate similar performance across all formats.
@@ Appendix D Appendix: Other Models
We experimented with the Llama-3.3 70B instruction-tuned model, as it was promoted to be more efficient with “enhanced” quality (cf. the official model card). However, our evaluation does not show any considerable improvement compared to Llama-3 70B (cf. Table 16), probably due to the different types of tasks they target.

### v4

@@ 5 Conclusions
After a few years of research on leveraging open-source LLMs for direct translation among other tasks, several businesses have been investigating the feasibility of integrating such general-purpose models into their specialized translation workflows. Among the main advantages of employing the in-context learning feature of LLMs is maintaining consistency, domain terminology and style throughout the entire document Dettmers et al. (2023); Wang et al. (2023). This can be achieved by feeding an LLM with similar translation pairs, preapproved terms, or other useful data at inference time, hoping that the LLM’s output will be influenced by the provided linguistic data. Moreover, several businesses and organizations have resorted to open-source LLMs for data privacy and security reasons. Nevertheless, deploying open-source LLMs into production poses efficiency challenges, including infrastructure requirements, latency at translation time, and inconsistent quality. This raises the question of whether conventional encoder-decoder MT models remain relevant, particularly for domain-specific translation.
As elaborated in Section 3, we evaluated a range of LLMs in the medical domain. Moreover, we fine-tuned selected models for zero-shot and one-shot translation. Our experiments covered both the multilingual encoder-decoder MT model, NLLB-200 3.3, and several autoregressive decoder-only LLMs, such as Gemma, Llama, and Mistral. Our retrieval-augmented translation used one-shot in-context examples based on semantic similarly with the new source (fuzzy match), which achieves better results than random selection (Agrawal et al., 2023; Moslem et al., 2023a). In the future, we would like to compare retrieval-augmented context (as in this work) with direct context from previous and next sentences in the document.
@@ 7 Limitations
Since all the data used in this work has been publicly available for years, there is a potential for data contamination. Many open-source LLMs do not disclose the specific datasets used for training, so we cannot be certain whether our test datasets were part of the training data for any of these models. However, we conducted multiple experiments for each model, including sentence-level translation (zero-shot), retrieval-augmented translation with fuzzy matches (one-shot), and fine-tuning of some models. As such, we consider our results a reliable indication of the translation quality that can be expected from these LLMs in similar setups. It is important to note that, as this paper demonstrates, results may vary across different languages or domains.
@@ A.1 Filtering
The filtering step removes duplicates, too long segments, and semantically mismatching segments. The scripts of our data preparation are publicly available. [^11] https://github.com/ymoslem/MT-Preparation/
Dataset Raw Filtered
ELRC_2922 4,365 4,022
ELRC_2923 502 470
ELRC_3382 6,593 6,537
ELRC-antibiotic 1,072 1,009
ELRC-EUROPARL_covid 789 747
ELRC-presscorner_covid 11,361 11,265
ELRC-wikipedia_health 4,366 4,045
EMEA 1,092,568 292,098
Tico-19 3,071 2,768
Total 1,124,687 322,961
@@
Dataset Raw Filtered
ELRC_2922 7,292 7,129
ELRC_2923 560 519
ELRC_3382 3,513 3,465
ELRC-antibiotic 839 775
ELRC-EUROPARL_covid 864 803
ELRC-presscorner_covid 6,319 6,239
ELRC-wikipedia_health 7,293 7,153
EMEA 1,082,144 290,782
SciELO 3,084,830 2,716,409
Tico-19 3,071 3,067
Total 4,196,725 3,036,341
@@
Domain Dataset Raw Filtered
Medical ELRC-wikipedia_health 608 593
Tico-19 3,071 3,056
Total 3,679 3,649
@@
Generic GlobalVoices 32,307 29,623
Wikimedia 16,289 13,829
Total 48,596 43,452
@@ A.2 Sampling and Splitting
A.3 LLM Prompts
Zero-shot prompt example:
  • English: <source_sentence>
  • Swahili:
One-shot prompt example:
  • English: <source_fuzzy_sentence>
  • Swahili: <target_fuzzy_sentence>
  • English: <source_sentence>
  • Swahili:
@@ B.2.2 NLLB
B.3 Dataset Packing
This section examines the performances of the models with and without packing. Packing dataset, initially introduced by (Raffel et al., 2020), helps increase the training efficiency by enabling fine-tuning on more sequences. In this method, multiple short examples are packed in the same input sequence. Packing tokens into the maximum sequence length helps improve training efficiency. While packing is primarily intended to improve the training throughput, our experiments demonstrate it also improves model performance. In our experiments, we use the HuggingFace’s SFTTrainer, which supports example packing.
@@ C.1 CTranslate2 vs vLLM
Table 14: Evaluation of the translation performance of NLLB-200 3.3B and Mistral 7B, without quantization and with lower-precision quantization formats (float16 and int8) using CTranslate2. The results indicate similar performance across all formats.
C.2 Quantization
Moreover, we compared generation with Llama-3 70B with and without Activation-aware Weight Quantization (AWQ) using vLLM. As Table 15 shows, results without quantization are slightly better. All the results in the main paper for LLMs are for non-quantized models, except Llama-3.1 405B and DeepSeek V3 685B, due to computing constrains. Hence, we tested Llama-3 70B with and without quantization and reported the results in this section to assess how AWQ affects the quality.
Lang Context Quant. BLEU ↑ ChrF++ ↑ TER ↓ COMET ↑
EN-FR zero-shot None 39.74 62.51 53.36 66.75
AWQ 39.02 61.97 54.69 64.97
one-shot None 51.93 70.57 43.03 76.76
AWQ 51.83 70.43 43.00 76.19
EN-PT zero-shot None 43.86 67.35 43.95 84.05
AWQ 43.12 66.59 45.58 82.12
one-shot None 45.95 68.82 41.68 86.58
AWQ 45.59 68.57 42.12 86.10
@@ Appendix D Appendix: Other Models
We evaluated the Llama-3.3 70B instruction-tuned model, as it was promoted to be more efficient with “enhanced” quality (cf. the official model card). However, our evaluation does not show any considerable improvement compared to Llama-3 70B (cf. Table 16), probably due to the different types of tasks they target.